A student mixed 100.0 mL of 1.50 mol/L sulfuric acid with 200.0 mL of 1.50 mol/L sodium hydroxide. Both solutions were at 19.670C initially and the highest temperature reached by the reaction mixture was 34.060C. Calculate the molar enthalpy (deltaH) of neutralization for sulfuric acid.
1
Expert's answer
2019-07-10T02:52:55-0400

Solution.

"Q(solution) = (m(acid) + m(hydroxide)) \times c(H2O) \times (t2-t1)""m(acid) = C(acid) \times V(acid) \times M(acid)""m(hydroxide) = C(hydroxide) \times V(hydroxide) \times M(hydroxide)"

Q(solution) = (14.7+12)*4.18*(34.06-19.67) = 1606.01 J


"Q(reaction) = -Q(solution)"

"\Delta H = \frac{Q(reaction)}{n(lim)}"

n(lim) is the limiting reagent in the reaction.

n(acid) = 0.15 mole

n(hydroxide) = 0.3 mole

Acid is limiting reagent.


"\Delta H = \frac{-1606.01}{0.15}"

"\Delta H = -10706.73 J"

Answer:

"\Delta H = -10706.73 J"